OpenStudy (anonymous):

Show all work. Find the following product: (y-1)^2

OpenStudy (anonymous):

first write \[(y-1)(y-1)\] then do 4 multiplications

jimthompson5910 (jim_thompson5910):

\[\large (y-1)^2\] \[\large (y)(y)+(y)(-1)+(-1)(y)+(-1)(-1)\] \[\large y^2-y-y+1\] \[\large y^2-2y+1\] So \[\large (y-1)^2=y^2-2y+1\]

OpenStudy (anonymous):

with a little practice you will see that \[(a+b)^2=a^2+2ab+b^2\] and \[(a-b)^2=a^2+2ab+b^2\]but for now it is probably best to write out all 4 multiplications and combine like terms as jimthompson did

OpenStudy (anonymous):

sorry \[(a-b)^2=a^2-2ab+b^2\]